One number is 17 less than another; their sum is 125. What are the numbers?
Answer by Cromlix(4381)

One number = x
Other number = x - 17
x + (x - 17) = 125
x + x - 17 = 125
2x = 142
x = 71
One number = 71
Other number = 54
Hope this helps.
:-)